Antonio Santos Facultad de Mateméticas Universidad de Vigo Spain E-mail address: asisanto@clix.pt

Banach Spaces and Polytopes

Let X be a Banach Space and B(X) the unit ball of X. If B(X) is a polyhedron then B(X^{∗}) is a polyhedron? Where B(X^{∗}) is a unit ball of the dual space.